# ODMA: On-Demand Memory Allocation Framework for LLM Serving on LPDDR-Class Accelerators

### v1

@@ Abstract
On Alpaca [10] and Google-NQ [18], ODMA improves S3’s predictor accuracy from 98.60% to 99.55% and from 82.68% to 93.36%, respectively (baseline numbers from [5]). Serving DeepSeek-R1-Distill-Qwen-7B on a node with four Cambricon MLU370-X4 accelerators (LPDDR5), ODMA raises device-memory utilization on Alpaca from 55.05% to 72.45%, and improves requests-per-second (RPS) and tokens-per-second (TPS) by 29% and 27% over a Cambricon-vLLM static pre-allocation baseline. These results demonstrate that predictor-driven, hardware-aware allocation unlocks efficient LLM serving on RACM platforms without hardware changes, complementary to paging-style designs such as PagedAttention on HBM systems [14].
@@ 1 Introduction
  • A Cambricon-vLLM-compatible prototype on LPDDR5-based MLU370 accelerators. On Alpaca and Google-NQ workloads, ODMA improves length-prediction accuracy over S3 [5], and increases device-memory utilization and end-to-end throughput relative to a static worst-case pre-allocation baseline.
@@ 2.2 Random-Access-Constrained Device Memory (RACM)
3 Problem Formulation
Let $L$ be the random generation length with (time-varying) distribution $F(\ell)$. Static allocators fix a worst-case $L_{\max}$ [5, 26], while paged allocators [14] amortize fragmentation via small blocks but rely on high random-access tolerance. We seek an LPDDR-aware on-device allocator with: (i) high utilization (low internal fragmentation), (ii) robustness to heavy tails and drift, (iii) preservation of sequential access, and (iv) minimal integration burden for vLLM-like stacks.
@@ 4.1 High-Level Architecture
Figure˜1 illustrates the overall serving pipeline. At a high level, ODMA inserts a memory-aware control path alongside an otherwise standard LLM serving stack:
  1. 1.
    Request arrival. User prompts arrive at the front-end and are forwarded to the Predictor.
  2. 2.
    Prediction and tagging. The Predictor estimates the generation length $\hat{L}$ for each request and attaches a bucket tag encoding the expected KV-cache size (and, when available, an uncertainty score).
  3. 3.
    Task Pool. Tagged requests are stored in a centralized Task Pool. The pool buffers outstanding requests and exposes their bucket tags to the Scheduler.
  4. 4.
    Scheduler. The Scheduler selects tasks from the Task Pool and forms inference batches. It prefers grouping tasks with compatible bucket tags, which simplifies allocation and reduces fragmentation.
  5. 5.
    Runtime. Once a batch is formed, the Runtime initiates decoding. Before the first token is generated, it requests KV-cache memory for each task from the Allocator according to its bucket tag.
  6. 6.
    Allocator and Memory Pool. The Allocator manages a global Memory Pool of device-memory blocks. For each request, it chooses a block with the appropriate bucket size from the pool on a target accelerator, ensuring a contiguous layout in physical memory. During decoding, the Runtime performs attention over these contiguous KV blocks, which translates into streaming-friendly accesses on LPDDR.
  7. 7.
    Completion and feedback. After the response is produced, the Runtime returns the KV block to the pool. The realized generation length is logged and fed back into an online histogram, which is used by the dynamic bucket manager to periodically refresh bucket boundaries.
Across multiple devices, a Cluster Manager inside the Allocator assigns batches to accelerator instances (e.g., multiple MLU/NPU cards) and coordinates their per-device Memory Pools. Each device maintains its own pool of blocks organized by bucket size (e.g., min/mid/max blocks), while the Cluster Manager ensures that load and memory pressure stay balanced.
This architecture cleanly separates concerns: the Predictor estimates per-request demand, the Task Pool and Scheduler shape workload mix and batch composition, and the Allocator enforces RACM-aware contiguous memory allocation across devices.
@@ 4.2 Generation-Length Predictor
The ODMA Predictor follows the spirit of S3 [5] but is integrated more tightly with the underlying memory system. For each request, the Predictor ingests:
  • tokenized prompt features (length, positional statistics, etc.),
  • lightweight metadata (request type, temperature, system prompt ID),
  • optional contextual signals (e.g., conversation depth).
These features are encoded by a compact Transformer or similar encoder, and the head network produces:
  1. 1.
    a point estimate $\hat{L}$ of the generation length, and
  2. 2.
    an uncertainty score $u$ (e.g., derived from a softmax over coarse length buckets or from the variance of an ensemble).
The value $\hat{L}$ is used to select a bucket, while $u$ determines whether the request should be treated as “risky” and routed directly to the large bucket (see Section˜4.4). Because predictor computation is performed once per request and amortized over potentially hundreds of decode steps, even a moderately sized model adds negligible overhead compared with LLM inference.
Predictor training uses offline traces (e.g., Alpaca [10], Google-NQ [18], and internal service logs). During deployment, fine-grained logging of realized lengths enables periodic re-training or lightweight online adaptation, which helps track distribution shifts common in production services.
Figure 2: Predictor pipeline. Prompt features and metadata are fed into a lightweight encoder to produce a length estimate $\hat{L}$ and uncertainty $u$. These outputs determine the bucket tag attached to each request.
@@ 4.3 Dynamic Bucket Manager
Static bucket boundaries, as used in S3 [5], implicitly assume a fixed generation-length distribution. Real-world traffic, however, is non-stationary: peaks in prompt length, changes in application mix, and new product features all shift the distribution of generation lengths. If bucket boundaries remain static under such drift, two problems arise:
  • Fragmentation. Many requests fall far below the upper bound of their bucket, wasting KV-cache memory.
  • Misprediction amplification. Buckets that no longer match the current distribution exacerbate the impact of prediction errors, increasing both under- and over-provisioning.
ODMA addresses this by maintaining an online histogram of realized lengths and periodically recomputing bucket boundaries. Let $\{L_{j}\}$ be the sequence of realized generation lengths over a sliding window. The system maintains empirical quantiles $Q_{p_{1}},\dots,Q_{p_{B}}$ and defines bucket upper bounds:
@@
The quantiles $p_{i}$ are chosen to roughly equalize bucket occupancy while respecting hardware constraints (e.g., maximum number of buckets, alignment requirements). When the window is refreshed, the bucket manager updates $\mathcal{B}$; subsequent allocations use the new boundaries, while existing allocations remain unchanged.
Given a prediction $\hat{L}$ and uncertainty-adjusted value $\hat{L}^{\prime}=\hat{L}\cdot(1+\alpha u)$ (with a small inflation factor $\alpha$), the allocator chooses the smallest bucket whose bound $b_{i}$ exceeds $\hat{L}^{\prime}$. This scheme:
  • keeps bucket sizes aligned with the live workload,
  • reduces internal fragmentation compared with static buckets,
  • preserves contiguous KV-cache layout (only a small number of coarse buckets are used, not fine-grained pages),
  • naturally co-adapts with the Predictor and Scheduler.
Because only bucket boundaries are updated, not the physical layout of existing blocks, boundary refresh is cheap and does not disrupt in-flight requests.
@@ 4.4 Large-Bucket Safety Strategy
Even with accurate prediction and adaptive buckets, heavy-tailed distributions produce occasional large generations. On HBM GPUs, paged designs can absorb such outliers using many small blocks with low penalty. On RACM accelerators, however, frequent random accesses to sparse blocks are expensive.
ODMA therefore introduces a large bucket that acts as an insurance pool for atypical or high-uncertainty requests. It serves two complementary roles:
High-uncertainty routing.
If the uncertainty score $u$ exceeds a configurable threshold $\tau$, ODMA directly assigns the request to the large bucket. This avoids repeated resizing for requests whose true length is difficult to anticipate.
Mid-decode overflow handling.
If a request allocated from a regular bucket grows beyond the bucket upper bound $b_{i}$ during decoding, the Runtime triggers an overflow procedure:
  1. 1.
    allocate a larger KV block from the large bucket,
  2. 2.
    perform a sequential copy of the existing KV-cache into the new block,
  3. 3.
    update internal pointers to reference the new block,
  4. 4.
    return the old block to the regular bucket pool.
Because KV-cache accesses are sequential, the copy itself is bandwidth-efficient on LPDDR. Moreover, overflows are rare in practice due to the combination of prediction, inflation by $u$, and quantile-based bucket design. The large bucket thus enables narrower regular buckets (and higher memory utilization) without sacrificing robustness.
@@ 4.5 Allocator and Runtime Integration
The Allocator is implemented as a thin layer between the Runtime and the underlying device memory. It exposes two main APIs:
  • reserve(request_id, batch_id): called when a request is admitted into a batch. It runs the Predictor (if not already run), consults the current bucket boundaries, and returns a pointer to a contiguous KV-cache region of the appropriate size.
  • release(request_id): called when the request finishes decoding. It returns the block to the free list of its bucket and records the realized generation length for histogram updates.
The global Memory Pool on each accelerator is partitioned into bucket-specific free lists. To match typical RACM hardware, block sizes are aligned to a small number of granularities (e.g., min, mid, max blocks), similar to the illustrative layout in Figure˜1. This keeps allocator metadata compact and lookups fast.
In multi-accelerator deployments, the Cluster Manager:
  • tracks free capacity and bucket occupancy on each device,
  • assigns incoming batches to devices to balance both compute and memory,
  • keeps the logical bucket partitioning synchronized across devices so that scheduling decisions remain stable.
A background maintenance thread periodically:
  • aggregates recent length samples into histograms,
  • recomputes bucket boundaries,
  • updates allocator configuration for future allocations.
All of these operations are off the critical path of token generation. Attention kernels and model implementations remain unchanged; ODMA interposes only at allocation sites, making it straightforward to integrate into vLLM-like frameworks and other production inference stacks.
@@ 6 Evaluation
We evaluate ODMA on Alpaca and Google-NQ [10, 18] using the hardware and software stack described in Section˜5. We report (i) length-prediction accuracy, comparing ODMA against the S3 predictor [5], and (ii) end-to-end throughput and device-memory utilization, comparing ODMA against a static worst-case pre-allocation baseline implemented in Cambricon-vLLM 0.6.2. Unless otherwise stated, all methods serve DeepSeek-R1-Distill-Qwen-7B with identical model weights and decoding parameters.
@@ 6.2 Throughput and Utilization
Figure˜4 reports end-to-end throughput relative to the static pre-allocation baseline. On the MLU370 node, ODMA improves RPS by 29% and TPS by 27% by reducing internal fragmentation and admitting more concurrent requests under the same device-memory budget.
Figure 4: Throughput improvement of ODMA over a static pre-allocation baseline (Cambricon-vLLM). Left: RPS; Right: TPS.
@@
7 Discussion and Limitations
Workload dependence.
ODMA adapts boundaries from observed lengths; cold starts or domain shifts may temporarily reduce utilization. Seeding with offline traces and conservative initial buckets mitigates this, similar to adaptive scheduling strategies discussed in recent surveys [12, 17].
Applicability.
ODMA targets RACM memory where random access is materially slower than streaming (e.g., LPDDR5-based accelerators such as Cambricon MLU370 [6, 2]). On HBM systems, paged designs [14] or hybrid approaches (e.g., decoupling prefill/decode [30, 31, 32]) may be preferable; integrating ODMA with such engines is future work.

### v2

@@ Abstract
On Alpaca [10] and Google-NQ [18], ODMA improves S3’s predictor accuracy from 98.60% to 99.55% and from 82.68% to 93.36%, respectively (baseline numbers from [5]). Serving DeepSeek-R1-Distill-Qwen-7B on a node with four Cambricon MLU370-X4 accelerators (LPDDR5), ODMA raises device-memory utilization on Alpaca from 55.05% to 72.45% and on Google-NQ from 42.54% to 61.79%, and improves throughput (TPS) by 23% on Alpaca and 27% on Google-NQ over a Cambricon-vLLM static pre-allocation baseline. These results demonstrate that predictor-driven, hardware-aware allocation unlocks efficient LLM serving on RACM platforms without hardware changes, complementary to paging-style designs such as PagedAttention on HBM systems [14].
@@ 1 Introduction
  • A Cambricon-vLLM-compatible prototype on LPDDR5-based MLU370 accelerators. On Alpaca and Google-NQ workloads, ODMA improves length-prediction accuracy over S3 [5], and increases device-memory utilization and end-to-end throughput (TPS) relative to a static worst-case pre-allocation baseline.
@@ 3.1 High-Level Architecture
Figure˜1 summarizes the end-to-end serving pipeline of ODMA. The key idea is to augment an otherwise standard LLM serving stack with a memory-aware control path that predicts per-request KV-cache demand, propagates this information through scheduling, and enforces RACM-aware contiguous allocation at runtime.
Requests enter the system at the front-end and are first processed by the Predictor, which estimates the generation length $\hat{L}$ and produces a bucket tag that encodes the expected KV-cache footprint (optionally augmented with an uncertainty score). The tagged requests are inserted into a centralized Task Pool that buffers outstanding work and exposes bucket tags as first-class scheduling signals. The Scheduler then selects tasks from the pool to form inference batches, preferentially co-batching requests with compatible bucket tags. This policy stabilizes per-batch memory demand and reduces allocation-induced fragmentation by aligning batch composition with the coarse bucket structure.
After batching, the Runtime begins decoding. Prior to generating the first token, it invokes the Allocator to reserve KV-cache space for each request according to its bucket tag. The Allocator manages a device-local Memory Pool organized as bucket-specific free lists and returns pointers to contiguous KV blocks in physical memory. During decoding, attention kernels operate over these contiguous regions, yielding streaming-friendly access patterns that match LPDDR-class memory systems and avoid the random-access penalties typical of paged KV designs on RACM accelerators. Upon request completion, the Runtime releases the corresponding KV block back to the pool. The realized generation length is logged and aggregated into an online histogram, closing the feedback loop used by the dynamic bucket manager to periodically refresh bucket boundaries and track workload drift.
In multi-accelerator deployments, ODMA extends this design with a Cluster Manager integrated into the Allocator. Each accelerator maintains its own bucketed Memory Pool (e.g., a small set of aligned block granularities such as min/mid/max), while the Cluster Manager assigns batches to devices by jointly considering compute load and memory pressure. It also keeps the logical bucket configuration synchronized across devices, ensuring that scheduling decisions remain stable as the system adapts.
@@ 3.2 Generation-Length Predictor
ODMA uses a lightweight generation-length predictor to estimate the KV-cache demand of each request before decoding starts, in the spirit of S3 [5] but more tightly coupled with our bucketed memory pool. Given an incoming prompt, the predictor extracts compact features from the tokenized input (e.g., prompt length and simple positional statistics) together with lightweight request metadata such as request type, decoding temperature, and system-prompt identifier; when available, it also incorporates contextual signals like conversation depth. These signals are designed to be cheap to compute and to remain stable across model backends, enabling a predictor that generalizes across serving configurations.
The features are encoded by a compact Transformer (or an equivalent small encoder), followed by a head that produces (i) a point estimate of the generation length, denoted by $\hat{L}$, and (ii) a scalar uncertainty score $u$ that reflects the predictor’s confidence. In practice, $u$ can be instantiated using the entropy or margin of a softmax distribution over coarse length buckets, or by the variance of a small ensemble. The allocator uses $\hat{L}$ to select the smallest bucket that can accommodate the predicted KV footprint, while $u$ governs conservatism: high-uncertainty requests are treated as risky and routed directly to the large bucket (see Section˜3.4), and moderate uncertainty can be handled by inflating the effective estimate (e.g., via an uncertainty-aware safety factor) to reduce under-allocation. This design explicitly prioritizes avoiding catastrophic underestimation—whose consequence is mid-decode overflow—while still keeping regular buckets tight for high utilization.
Predictor inference is executed once per request and is amortized over hundreds of decode steps, so its overhead is negligible relative to LLM decoding. We train the predictor offline using representative traces (Alpaca [10], Google-NQ [18], and internal service logs), and we continually log realized generation lengths in deployment to support periodic re-training or lightweight online calibration. This feedback loop helps the predictor track the distribution shifts that commonly occur in production workloads.
Figure 2: Generation-length prediction in ODMA. A lightweight encoder maps prompt features and request metadata to a length estimate $\hat{L}$ and an uncertainty score $u$, which together determine the bucket tag used by the allocator.
@@ 3.3 Dynamic Bucket Manager
Static bucket boundaries (e.g., S3 [5]) effectively assume that the generation-length distribution is stable. In production, however, request traffic is inherently non-stationary: prompt lengths exhibit diurnal spikes, the application mix evolves, and new features can introduce long-tail behaviors. Under such drift, fixed boundaries gradually decouple from the live workload, leading to increased internal fragmentation (requests under-utilize their assigned bucket capacity) and to amplified consequences of prediction errors (buckets become either too coarse or poorly placed, increasing both over- and under-provisioning).
To remain workload-aligned, ODMA maintains an online histogram of realized generation lengths and periodically refreshes bucket boundaries using a sliding window of recent requests. Let $\{L_{j}\}$ denote realized generation lengths observed within the current window. The manager estimates empirical quantiles $Q_{p_{1}},\dots,Q_{p_{B}}$ and sets the bucket upper bounds as
@@
The quantile levels $\{p_{i}\}$ are selected to roughly equalize bucket occupancy, subject to hardware and implementation constraints such as the maximum number of buckets and alignment requirements. This yields a small set of coarse buckets that track the evolving workload without resorting to fine-grained paging.
When the window advances, the bucket manager updates $\mathcal{B}$ and the allocator applies the new boundaries only to subsequent requests; in-flight requests keep their originally assigned blocks. This update semantics avoids repacking or migrating KV-cache memory, ensuring that boundary refresh is cheap and does not disrupt ongoing decoding.
At allocation time, the predictor provides a length estimate $\hat{L}$ and an uncertainty score $u$. We incorporate conservatism via an uncertainty-aware inflation,
$\hat{L}^{\prime}=\hat{L}\cdot(1+\alpha u),$
where $\alpha$ is a small tuning factor. The allocator then chooses the smallest bucket whose upper bound satisfies $b_{i}\geq\hat{L}^{\prime}$. By continuously adapting $\mathcal{B}$ to recent realizations while using $\hat{L}^{\prime}$ to hedge against uncertainty, ODMA simultaneously reduces internal fragmentation, limits the damage of misprediction under distribution shift, and preserves a contiguous KV-cache layout with only a small number of coarse buckets. Moreover, because the refreshed buckets reflect the live traffic, they naturally co-adapt with the Predictor and the Scheduler, improving end-to-end stability under changing workloads.
@@ 3.4 Large-Bucket Safety Strategy
Even with an accurate predictor and workload-adaptive bucket boundaries, the generation-length distribution in real deployments is typically heavy-tailed, and sporadic long generations are inevitable. On HBM-based GPUs, paged KV-cache designs can accommodate such outliers by chaining many small blocks, often with acceptable overhead. In contrast, on RACM accelerators backed by LPDDR, this approach is substantially less attractive: assembling sparse blocks induces frequent non-contiguous memory accesses, which are costly and can dominate decode-time latency.
To provide robustness without adopting fine-grained paging, ODMA introduces a dedicated large bucket that acts as an insurance pool for atypical requests. Conceptually, the large bucket complements the regular buckets by handling cases where (i) the predictor is uncertain, or (ii) the realized generation exceeds the initially provisioned capacity. This strategy enables regular buckets to remain relatively narrow—improving utilization—while still guaranteeing correctness under long-tail events.
Uncertainty-aware routing. For each request, the predictor outputs an uncertainty score $u$. When $u$ exceeds a configurable threshold $\tau$, ODMA bypasses regular buckets and allocates directly from the large bucket. This conservative routing avoids repeated re-allocation for requests whose realized length is intrinsically hard to anticipate (e.g., due to ambiguous prompts or rapidly shifting workloads), thereby stabilizing both latency and allocator behavior.
Mid-decode overflow recovery. A second entry point arises when a request initially placed in a regular bucket outgrows its bucket upper bound $b_{i}$ during decoding. In this case, the runtime invokes an overflow procedure that migrates the request into the large bucket: it allocates a larger KV block from the large bucket, performs a sequential copy of the existing KV-cache into the new block, atomically updates internal references to the new block, and finally returns the old block to the original regular-bucket pool. Although migration introduces additional work, the copy is bandwidth-efficient on LPDDR because it is a contiguous, streaming transfer rather than a sequence of random gathers. Furthermore, overflows are rare in practice due to the combined effect of length prediction, uncertainty-aware inflation, and quantile-based bucket placement, so the amortized overhead is small.
@@ 3.5 Allocator and Runtime Integration
ODMA implements the Allocator as a thin mediation layer between the Runtime and device memory, providing a minimal interface that can be hooked into vLLM-like inference stacks without modifying attention kernels or model code. Concretely, the Allocator exposes two entry points. The first, reserve(request_id), is invoked when a request is admitted into a batch: it runs the Predictor if the request has not been annotated yet, queries the current bucket boundaries (including the large-bucket policy), and returns a pointer to a contiguous KV-cache region sized according to the selected bucket. The second, release(request_id), is called when the request finishes decoding: it returns the corresponding block to the free list of its bucket and logs the realized generation length, which is later consumed by the dynamic bucket manager to refresh its online histogram.
On each accelerator, the Allocator manages a global Memory Pool that is logically partitioned into bucket-specific free lists. To match typical RACM hardware and keep metadata compact, block sizes are restricted to a small set of aligned granularities (e.g., min/mid/max blocks, consistent with Figure˜1); this design yields constant-time lookups and avoids the bookkeeping overhead associated with fine-grained paging while preserving the streaming-friendly contiguous layout required by LPDDR-class memories.
In multi-accelerator deployments, the Allocator further embeds a Cluster Manager that coordinates per-device pools. The Cluster Manager continuously tracks device-local free capacity and bucket occupancy, assigns incoming batches to devices to balance both compute and memory pressure, and keeps the logical bucket partitioning synchronized across devices so that scheduling and placement decisions remain stable under drift.
Finally, ODMA runs a lightweight maintenance routine off the decode critical path: a background thread periodically aggregates recent length samples into histograms, recomputes bucket boundaries, and updates allocator configuration for subsequent allocations (leaving in-flight requests untouched). As a result, all adaptive behavior is confined to the control plane and allocation sites, while the token-generation inner loop (including attention kernels) remains unchanged, making integration straightforward in production inference stacks.
@@ 5 Evaluation
We evaluate ODMA on Alpaca and Google-NQ [10, 18] using the hardware and software stack described in Section˜4. We report (i) length-prediction accuracy, comparing ODMA against the S3 predictor [5], and (ii) end-to-end throughput and device-memory utilization, comparing ODMA against a static worst-case pre-allocation baseline implemented in Cambricon-vLLM 0.6.2. Unless otherwise stated, all methods serve DeepSeek-R1-Distill-Qwen-7B with identical model weights and decoding parameters.
@@ 5.2 Throughput and Utilization
Figure˜4 reports end-to-end throughput measured in tokens per second (TPS) relative to the static pre-allocation baseline. On the MLU370 node, ODMA improves TPS by 23% on Alpaca and 27% on Google-NQ by reducing internal fragmentation and admitting more concurrent requests under the same device-memory budget.
Figure 4: Throughput (TPS) improvement of ODMA over a static pre-allocation baseline (Cambricon-vLLM). Left: Alpaca; Right: Google-NQ.